# RAG-DDR: Optimizing Retrieval-Augmented Generation Using Differentiable Data Rewards

### v1

@@ 6 Conclusion
This paper proposes Differentiable Data Rewards (DDR), a method aimed at end-to-end optimizing the Retrieval-Augmented Generation (RAG) model using the DPO method. DDR optimizes each agent by collecting the reward in a rollout way and aligns data preferences among these communicative agents. We build a two-agent RAG system and optimize it using DDR to implement the RAG-DDR model. Our experiments demonstrate that DDR helps the generation module produce responses of an appropriate length and avoids overfitting the training signals during SFT. Our further analyses reveal that the DDR optimized generation model can better capture key information from retrieved documents and mitigate the conflict between external knowledge and parametric memory.
@@ A.1 Additional Experimental Details
In this subsection, we first outline the process of constructing the training data. Then we show the prompt templates used in our experiments. In experiments, we keep the same training data, hyperparameters, and epochs to implement both RA-DIT and RAG-DDR for a fair comparison.
@@
Figure 4: Prompts of Different Training and Evaluating Tasks.
Figure 5: Prompts Used for Knowledge Refinement.
@@
A.2 Additional Ablation Studies on RAG-DDR
@@
A.3 Characteristics of The Knowledge Refinement Module of RAG-DDR
As shown in Figure 6, we explore the characteristics of the knowledge refinement module ($V_{\text{KR}}$) optimized using different methods, including Vanilla RAG, RA-DIT and RAG-DDR.
[…]
As shown in Figure 6(a), we calculate the accuracy of top-5 documents by Vanilla RAG, RA-DIT, and RAG-DDR. The retrieval accuracy evaluates whether or not the retained top-5 documents contain the ground truth. If $V_{\text{KR}}$ discards all the documents retrieved, the accuracy is 0. RAG-DDR outperforms Vanilla RAG and RA-DIT on all tasks. It indicates that the DDR method can make $V_{\text{KR}}$ accurately retain documents that contain the necessary knowledge to answer the query. As shown in Figure 6(b), we use the refined documents by different knowledge refinement module ($V_{\text{KR}}$) to augment the DDR trained generation module ($V_{\text{Gen}}$). DDR-RAG also outperforms other models, indicating that DDR can help better align data preferences between $V_{\text{Gen}}$ and $V_{\text{KR}}$.
@@ A.3 Characteristics of The Knowledge Refinement Module of RAG-DDR
A.4 The General LLM Ability of DDR Optimized Generation Module

### v2

@@ 6 Conclusion
This paper proposes Differentiable Data Rewards (DDR), aiming at end-to-end optimizing the Retrieval-Augmented Generation (RAG) model using the DPO method. DDR optimizes each agent by collecting the reward in a rollout way and aligns data preferences among these communicative agents. We build a two-agent RAG system and optimize it using DDR to implement the RAG-DDR model. Our experiments demonstrate that DDR helps the generation module produce responses of an appropriate length and avoids overfitting the training signals during SFT. Our further analyses reveal that the DDR optimized generation model can better capture key information from retrieved documents and mitigate the conflict between external knowledge and parametric memory.
Acknowledgments
This work is partly supported by the Natural Science Foundation of China under Grant (No. 62206042, No. 62137001, and No. 62272093), CCF-zhipu Large Model Innovation Fund (No. 202403).
@@ A.1 Additional Experimental Details
In this subsection, we first outline the process of constructing the training data. Then we show the prompt templates used in our experiments. Then, we describe the maximum generation length of RAG-DDR during inference. Finally, we provide details on the implementation of RAG-DDR training. To ensure a fair comparison, we train both RA-DIT and RAG-DDR using the same dataset, hyperparameters, and number of epochs. For Self-RAG, we adhere to it respective training data formats and hyperparameter settings.
@@
Maximum Generation Length. For the evaluation dataset, we set the maximum generation length for MARCO QA task to 100, while the maximum generation length for the rest tasks is set to 32.
@@
Figure 4: Prompt Templates of Different Training and Evaluating Tasks.
Figure 5: Prompt Templates Used for Knowledge Refinement.
@@
Training Details. For DDR training, we use automatic metrics such as Rouge-L and Accuracy to calculate the reward and set $\beta=0.1$. The learning rate is set to 5e-5, and each model is trained for one epoch. For the generation module, we feed 5 retrieved passages as external knowledge for augmenting the generation process. To optimize both the knowledge refinement module and generation module, we use LoRA (Hu et al., 2022) for efficient training.
Table 6: Overall Performance of RAG-DDR and Additional Baselines. The evaluation scores of InstructRAG on the MARCO QA and WoW tasks are usually low. The reason may lie in that the responses generated by InstructRAG are long and contain complex reasoning processes, making it difficult to conduct a fair evaluation for these string matching based evaluation metrics.
Method Open-Domain QA Multi-Hop QA Slot Filling Dialogue
NQ TriviaQA MARCO QA HotpotQA T-REx WoW
Short Generation Length
Filco (2023) 46.6 81.9 14.6 25.3 44.1 14.4
InstructRAG (2024) 4.2 21.1 14.3 11.0 17.1 9.2
RAG-DDR (w/ 2-Round) 52.1 89.6 27.3 39.0 40.6 16.8
Long Generation Length
InstructRAG (2024) 60.8 92.0 12.4 42.3 43.1 5.3
RAG-DDR (w/ 2-Round) 58.0 91.9 26.1 43.9 45.6 10.9
A.2 Additional Baseline Comparison Results
This section presents the comparison results between RAG-DDR (w/ 2-Round) and other baseline models. We employ InstructRAG (Wei et al., 2024) and Filco (Wang et al., 2023) as baseline models. InstructRAG guides the LLM to self-synthesize denoising instruction tuning data and trains the generation module to denoise the retrieved documents. Filco finetunes Flan-T5-XL (Chung et al., 2024) as a context filtering module to refine the retrieved documents. Specifically, the responses generated by InstructRAG include long reasoning and analysis, thus InstructRAG sets the maximum generation length as 4,096. For a fair comparison, we compare the performance of RAG-DDR and InstructRAG on the two maximum generation length settings: the short generation length (MARCO QA task is 100 and the rest tasks are 32), and the Long generation length (4,096). In our experiments, we employ Llama3-8B-Instruct as the knowledge refinement and generation modules.
As shown in Table 6, RAG-DDR outperforms Filco in almost all tasks. This suggests that the SFT method leads modules in the RAG system to overfit the training signals, making it less effective to optimize the entire RAG system to generate accurate responses. In contrast, in the long generation length setting, InstructRAG slightly outperforms RAG-DDR on NQ and TriviaQA tasks but performs slightly worse than RAG-DDR on T-REX and HotpotQA tasks. This illustrates that high-quality synthetic fine-tuning data can effectively optimize the performance of the generation module. However, in the short generation length setting, InstructRAG performs significantly worse than RAG-DDR, which indicates that compared to InstructRAG, RAG-DDR can generate accurate answers with shorter responses, without complex and long reasoning processes.
@@ A.2 Additional Baseline Comparison Results
A.3 Additional Ablation Studies on RAG-DDR
@@
Table 8: Effectiveness of RAG-DDR with More Agents. Llama3-8B is used as the backbone model for both the knowledge refinement and summarization modules, while MiniCPM-2.4B is used as the backbone model for the generation module.
Method Open-Domain QA Multi-Hop QA Slot Filling Dialogue
NQ TriviaQA MARCO QA HotpotQA T-REx WoW
Vanilla RAG (w/o $V_{\text{Sum}}$) 42.2 79.5 16.7 26.7 22.1 14.4
Vanilla RAG 44.1 82.3 17.5 28.9 25.4 15.4
RAG-DDR (Only Training $V_{\text{Gen}}$) 46.6 83.6 26.9 32.1 28.3 16.9
RAG-DDR (Training $V_{\text{Gen}}$ $V_{\text{Sum}}$) 47.6 83.9 26.4 32.0 28.6 17.0
RAG-DDR (All Training) 47.9 84.8 26.4 32.9 28.7 16.9
A.4 Effectiveness of RAG-DDR with More Agents
In this experiment, we extend RAG-DDR to more agents to evaluate its effectiveness in optimizing complex RAG systems. Specifically, we utilize the knowledge refinement ($V_{\text{KR}}$), summarization ($V_{\text{Sum}}$), and generation ($V_{\text{Gen}}$) modules to construct a new RAG system. In this system, the knowledge refinement module refines the retrieved documents $D$ to get the refined documents $\tilde{D}$, the summarization module summarizes the refined documents $\tilde{D}$ into a concise summary $\tilde{S}$ based on the query $q$, and the generation module uses this summary $\tilde{S}$ to answer the query $q$. This system can be represented as a three-agent system:
$\{q,D\}\rightsquigarrow V_{\text{KR}}\xrightarrow{\{q,\tilde{D}\}}V_{\text{Sum}}\xrightarrow{\{q,\tilde{S}\}}V_{\text{Gen}}\rightsquigarrow y_{\text{Gen}},$ (9)
where $\tilde{D}\subseteq D$. $\{q,D\}\rightsquigarrow$ and $\rightsquigarrow y_{T}$ represent sending the input $\{q,D\}$ to the RAG system and getting the final output $y_{T}$. In our implementation, we use Llama3-8B-Instruct as the backbone model to construct the knowledge refinement and summarization modules, and Minicpm-2.4B-sft as the backbone model to construct the generation module.
As shown in Table 8, we employ five models: Vanilla RAG (w/o $V_{\text{Sum}}$, Vanilla RAG, RAG-DDR (Only Training $V_{\text{Gen}}$), RAG-DDR (Training $V_{\text{Gen}}$ $V_{\text{Sum}}$) and RAG-DDR (All Training) to evaluate the effectiveness of the DDR method in RAG systems with more agents. RAG-DDR (Only Training $V_{\text{Gen}}$) indicates that we tune the Vanilla RAG using DDR by only optimizing the generation module ($V_{\text{Gen}}$). RAG-DDR (Training $V_{\text{Gen}}$ $V_{\text{Sum}}$) indicates that we use DDR to optimize both the generation module ($V_{\text{Gen}}$) and summarization module ($V_{\text{Sum}}$) in Vanilla RAG. RAG-DDR (All Training) indicates that we optimize all three modules in the Vanilla RAG.
Compared to Vanilla RAG (w/o $V_{\text{Sum}}$), Vanilla RAG demonstrates performance improvements across all evaluation tasks, showing the effectiveness of the summarization module in extracting relevant information from multiple documents. In contrast, RAG-DDR (Only Training $V_{\text{Gen}}$) shows greater improvements than Vanilla RAG, indicating that the primary improvements of RAG-DDR come from optimizing the generation module ($V_{\text{Gen}}$). When we subsequently optimize the summarization and knowledge refinement module based on RAG-DDR (Only Training $V_{\text{Gen}}$), the performance is further improved. This indicates that DDR not only improves the performance of the RAG system but also exhibits strong scalability, allowing it to be extended to different RAG systems.
Table 9: Effectiveness of RAG-DDR Using Large-Scale LLMs. In our experiments, we utilize Qwen2.5-14B-Instruct as the generation module and Llama3-8B-Instruct as the refinement module.
Method Open-Domain QA Multi-Hop QA Slot Filling Dialogue
NQ TriviaQA MARCO QA HotpotQA T-REx WoW
Vanilla RAG 48.1 81.1 18.5 29.4 37.3 14.3
RAG-DDR 51.2 84.1 23.0 35.1 44.6 16.6
A.5 Extending RAG-DDR to Larger-Scale LLMs
In this section, we explore the effectiveness of the DDR method in the RAG system that employs a larger-scale LLM as the backbone model of the generation module. Specifically, we utilize Qwen2.5-14B-Instruct (Yang et al., 2024) as the generation module and Llama3-8B-Instruct as the refinement module to build a new combination for the RAG system. As shown in Table 9, compared to vanilla RAG, RAG-DDR shows significant improvements across multiple tasks, highlighting the effectiveness of the DDR method in the RAG system with the larger-scale LLM.
@@ A.5 Extending RAG-DDR to Larger-Scale LLMs
A.6 Characteristics of the Knowledge Refinement Module of RAG-DDR
As shown in Figure 6, we explore the characteristics of the knowledge refinement module ($V_{\text{KR}}$) optimized using different methods, including zero-shot (Vanilla RAG), the SFT method (RA-DIT), and DDR (RAG-DDR).
[…]
As shown in Figure 6(a), we calculate the accuracy of top-$5$ documents by Vanilla RAG, RA-DIT, and RAG-DDR. The retrieval accuracy evaluates whether or not the retained top-5 documents contain the ground truth. If $V_{\text{KR}}$ module discards all the documents retrieved, the accuracy is 0. RAG-DDR outperforms Vanilla RAG and RA-DIT on all tasks. It indicates that the DDR method can make $V_{\text{KR}}$ module accurately retain documents that contain the necessary knowledge to answer the query. As shown in Figure 6(b), we use the refined documents by different knowledge refinement module ($V_{\text{KR}}$) to augment the DDR trained generation module ($V_{\text{Gen}}$). DDR-RAG also outperforms other models, indicating that DDR can help better align data preferences between $V_{\text{Gen}}$ and $V_{\text{KR}}$ modules.
A.7 The General LLM Ability of DDR Optimized Generation Module
@@
Table 10: The Effectiveness of Generation Module Optimized with Different Training Strategies.
Method Open-Domain QA Multi-Hop QA Slot Filling Dialogue
NQ TriviaQA MARCO QA HotpotQA T-REx WoW
MiniCPM-2.4B
Vanilla RAG 42.2 79.5 16.7 26.7 22.1 14.4
RA-DIT (Positive Label) 40.9 77.2 19.9 22.0 25.3 14.7
RA-DIT (Ground Truth) 41.8 78.6 19.6 26.1 25.2 15.7
RAG-DDR 46.8 81.7 28.3 31.2 32.2 17.0
Llama3-8B
Vanilla RAG 46.2 84.0 20.6 30.1 26.9 12.5
RA-DIT (Positive Label) 45.8 86.4 23.0 31.4 25.3 13.2
RA-DIT (Ground Truth) 46.2 87.4 20.3 34.9 41.7 14.8
RAG-DDR 50.2 87.8 25.2 36.9 36.2 14.8
Table 11: The Average Length of Responses Generated by Generation Module Optimized with Different Training Strategies.
Method Open-Domain QA Multi-Hop QA Slot Filling Dialogue
NQ TriviaQA MARCO QA HotpotQA T-REx WoW
MiniCPM-2.4B
Vanilla RAG 12.5 4.7 32.4 5.6 4.4 48.5
RA-DIT (Positive Label) 19.2 8.3 34.6 13.7 6.7 32.8
RA-DIT (Ground Truth) 7.6 4.8 13.8 8.9 5.0 17.2
RAG-DDR 27.4 16.9 47.8 27.2 11.2 56.6
Llama3-8B
Vanilla RAG 49.6 25.0 85.4 43.9 34.4 98.5
RA-DIT (Positive Label) 42.1 14.8 72.2 24.0 3.9 72.2
RA-DIT (Ground Truth) 6.9 8.7 14.5 23.6 4.0 16.7
RAG-DDR 56.0 33.9 81.2 51.6 35.2 101.5
A.8 The impact of different training strategies on the generation module
In this experiment, we further investigate the impact of different training strategies on the generation module $V_{\text{Gen}}$’s ability to utilize external knowledge and the average length of responses generated by $V_{\text{Gen}}$ module. We compare four models in this experiment, including Vanilla RAG, RA-DIT (Positive Label), RA-DIT (Ground Truth) and RAG-DDR. For RA-DIT (Positive Label), we employ the RA-DIT method to train the Vanilla RAG and regard the positive response in DDR training data as ground truth. For RA-DIT (Ground Truth), we employ the RA-DIT method to train the Vanilla RAG with the annotated labels.
As shown in Table 10, RAG-DDR consistently outperforms RA-DIT (Positive Label) and RA-DIT (Ground Truth) across different datasets, highlighting the effectiveness of the DDR method to enhance the ability of $V_{\text{Gen}}$ module to utilize external knowledge. In contrast, RA-DIT (Positive Label) performs worse than RA-DIT (Ground Truth), demonstrating that relying solely on positive samples for supervised fine-tuning is insufficient for effectively training LLMs.
Furthermore, we analyze the average length of the responses generated by $V_{\text{Gen}}$ module in Table 11. Compared to the RA-DIT (Ground Truth), RAG-DDR shows a more similar generation length distribution to the Vanilla RAG model, making the LLMs generate responses with a more appropriate length to answer the question. Notably, the response length of RA-DIT (Positive Label) is longer than RA-DIT (Ground Truth). This observation further shows that the SFT training method may lead LLMs to overfitting the training signals, which affects the response length of LLMs.